# A multi-messenger search for exotic field emission with a global magnetometer network

### v1

@@ Abstract
We present an analysis method to search for exotic low-mass field (ELF) bursts generated during large energy astrophysical events such as supernovae, binary black hole or binary neutron star mergers, and fast radio bursts using the Global Network of Optical Magnetometers for Exotic physics searches (GNOME). In our model, the associated gravitational waves or electromagnetic signals herald the arrival of the ELF burst that interacts via coupling to the spin of fermions in the magnetometers. This enables GNOME to serve as a tool for multi-messenger astronomy. The algorithm employs a model-agnostic excess-power method to identify network-wide candidate events to be subjected to a model-dependent generalized likelihood-ratio test to determine their statistical significance. We perform the first search with this technique on GNOME data coincident with the binary black hole merger S200311bg detected by LIGO/Virgo on the 11th of March 2020 and find no significant events. We place the first lab-based limits on combinations of ELF production and coupling parameters.
@@ I Introduction
Many extensions of the Standard Model posit the existence of new scalar or pseudoscalar fields with light ($m\ll 1\,\rm{keV/c^{2}}$) quanta, such as axions and axion-like particles (ALPs) PRESKILL1983127 ; ABBOTT1983133 ; DINE1983137 . In this work, we describe a method to search for light, ultra-relativistic scalar fields triggered by an astrophysical event using the Global Network of Optical Magnetometers for Exotic physics searches (GNOME). Following Dailey2021 , we refer to these fields as exotic low-mass fields (ELFs). ELFs are postulated to be generated during large energy astrophysical events such as supernovae Raf99 , binary black hole (BBH) or binary neutron star mergers baumann2019probing , and fast radio bursts (FRB) Tka15 . The detection of associated gravitational waves (GW) or electromagnetic (EM) signals herald the ELF signal arrival on Earth, and also indicates the GW/EM and ELF source location in the sky. By searching for signatures of the emissions from known sources shortly after a GW or EM trigger, we set limits on possible ELF production and coupling parameters associated with the observed astrophysical event.
@@
The production mechanism of ELFs is an open question arvanitaki2015discovering ; arvanitaki2017black ; baryakhtar2017black ; baumann2019probing ; baryakhtar2021black . Black holes may be surrounded by clouds of exotic bosons, with up to 10% of the black hole mass in the clouds. Theories of scalar-tensor gravity Fujii:2003pa describe cases of black holes and neutron stars being immersed in scalar fields. Such scenarios suggest that modes of these scalar fields may be excited during BBH or binary neutron star mergers franciolini2019effective . To set limits on ELF properties, we proceed with the following generic assumptions: (1) Some high-energy astrophysical event emits a spherical wave packet of the ELF $\phi(r)$ from a known location at distance $R$ from Earth. (2) The emitted wave packet carries away total energy $\Delta E$, where $\Delta E/\varepsilon_{0}\gg 1$ with $\varepsilon_{0}$ being the energy of an ELF quanta. Given the assumptions above, the ELF will have high mode occupation number and hence they are treated as classical phase-coherent waves.
@@
This manuscript is organized as follows. In Sec. II, we recall the characteristics of an ELF pulse, its propagation, and its coupling to atomic magnetometers. In Sec. III, we describe the GNOME network. In Sec. IV, we present our two-stage analysis pipeline. First, we look for large excess power in the network at particular times and frequencies. We then perform likelihood-ratio tests using the characteristics of the individual stations to determine the significance of possible events compared to time-shifted data, where we expect no signals. We perform the same test on injected signals to calibrate the network response to a known ELF signal. Finally, in Sec. V, we present results for one search target: the BBH merger S200311bg detected by LIGO/Virgo GCN_BBH .
@@ II.2 ELF detection with magnetometers
[…]
@@
with $\bf{S}$ being the atomic or nuclear spin Afach2023 . The ELF can be treated as a so-called “pseudomagnetic field” since it causes energy shifts of Zeeman sublevels in a manner similar to the effect of a real magnetic field. To explicitly make the analogy between the Zeeman Hamiltonian ($H_{Z}=-\boldsymbol{\mu}\cdot\boldsymbol{B}$, where $\boldsymbol{\mu}=\mu\boldsymbol{S}$ is the magnetic moment and $\boldsymbol{B}$ is a magnetic field) and the Hamiltonians describing ELF-spin interactions, we introduce the corresponding pseudomagnetic fields:
@@ III GNOME Network Characteristics
GNOME comprises a global network of magnetically shielded alkali-atomic magnetometers pustelny2013global ; afach2018characterization ; Afach2021 ; Afach2023 . These are located in over a dozen stations distributed over four continents. The sensors measure the projection of the ambient magnetic field on one sensitive axis via the variation of the Larmor spin precession of optically polarized alkali atoms in a vapor cell. Measurements are time-stamped using the Global Position System allowing for the use of coincident detection techniques to suppress unwanted background noises Afach2021 . In addition, a multi-layer mu-metal magnetic shield, present in each station, suppresses background magnetic fields by a factor of $\gtrsim 10^{6}$. Because GNOME magnetometers use atoms whose nuclei have a valence proton, they are sensitive to proton-spin couplings with the ELF. The expected ELF signal amplitudes measured by the sensors will be proportional to the relative contribution of the proton spin to the nuclear spin Afach2021 . ELFs can also couple to electron spins but the signal is attenuated by the electrons in the magnetic shields. To first order it was expected this attenuation is the same as the magnetic shielding factor kimball2016magnetic but the orbital angular momentum of the electrons in the shields can lessen this effect futurePaper . The coupling to electron spins is ignored in this manuscript.
@@
Station       Longitude Latitude       Az Alt $\sigma/g_{f}$
Beersheba 34.8043° E 31.2612° N 180°   $0.50^{+0.00}_{-0.19}$
Beijing 116.1868° E 40.2457° N 251° $-0.39^{+0.19}_{-0.00}$
Berkeley 122.2570° W 37.8723° N 90° $-0.39^{+0.19}_{-0.00}$
Canberra 149.1185° E 35.2745° S 90° $\ \ 0.70^{+0.00}_{-0.15}$
Daejeon 127.3987° E 36.3909° N 90° $-0.39^{+0.19}_{-0.00}$
Hayward* 122.0539° W 37.6564° N 90° $\ \ 0.70^{+0.00}_{-0.15}$
Hefei 117.2526° E 31.8429° N 90° $-0.38^{+0.05}_{-0.00}$
Krakow 19.9048° E 50.0289° N 45° $\ \ 0.50^{+0.00}_{-0.11}$
Lewisburg* 76.8825° W 40.9557° N 90° $\ \ 0.70^{+0.00}_{-0.15}$
Los Angeles* 118.4407° W 34.0705° N 270° $\ \ 0.50^{+0.00}_{-0.07}$
Mainz 8.2354° E 49.9915° N -90° $\ \ 0.50^{+0.00}_{-0.11}$
Moxa* 11.6147° E 50.6450° N 270° $-0.39^{+0.19}_{-0.00}$
Oberlin* 81.7796° W 41.2950° N 276° $-0.49^{+0.18}_{-0.00}$
Table 1: Location, orientation and proton spin couplings of sensors in GNOME. * marking denotes the stations that were active with quality data during the S200311bg event. (Moved from Fribourg to Moxa in 2018.)
For this work, each sensor has bandwidth of approximately 100 Hz. The response and sensitivity of each sensor is determined by calibration signals at multiple frequencies. Magnetic-field measurements are recorded at a sampling rate of 512 samples per second. In addition, a 1 sample per second data glitch / data quality indicator is also written to file. This boolean variable is derived from auxiliary sensors which monitor temperature, accelerations, magnetic field outside the shielding, rotation and other user-defined variables such as laser intensity, frequency, etc. If any of those quantities vary outside an allowed range within a given second, the data-quality flag – which is normally true – is switched to false. In our analysis, we disregard all the data having data quality flag false.
@@ V Search Target and Results
We now apply this analysis method to the BBH merger S200311bg (denoted as GW200311_115853 in Ref. PhysRevX.13.041039 ) detected by LIGO/Virgo on March 11th, 2020 at 11:58:53.398 UTC PhysRevX.13.041039 ; GCN_BBH . The luminosity distance is estimated to be $D_{L}=1.17^{+0.28}_{-0.40}$ Gpc at right ascension 00h08m and declination -07d27m with a sky localization of 35 deg2. The total mass before the merger was $61.9^{+5.3}_{-4.2}\,M_{\odot}$ and the final mass was $59.0^{+4.8}_{-3.9}\,M_{\odot}$.
@@ V.1 Foreground analysis
Figure 9 shows the cumulative number of foreground events above a given FPP versus the mean number of expected events. The foreground and time-shifted data are largely consistent. The most significant event had an expected mean of $\mu_{tot}=0.23$ or FPP$=0.21$, which would correspond to a detection with a significance of only $1.3\sigma$. While this is not significant, we calculate the corresponding ELF masses and energy scales as an example. This tile occurred at a delay time of 15920 s and a central frequency of 95.8 Hz. Through the FC method, we obtain a 90%-confidence upper limit of $\mathcal{B}_{\text{p}}<0.08$ pT. For the linear interaction, this corresponds to a mass of $m=2.1\times 10^{-19}$ eV [Eq. (33)] and linear energy scale of $f_{l}=87.4$ GeV at $\tau_{0}=1$ s and $\Delta E=1\,M_{\odot}c^{2}$ [Eq. (15)]. For the quadratic interaction, the mass is $m=1.0\times 10^{-19}$ eV and the quadratic energy scale is $f_{q}=0.2$ GeV at $\tau_{0}=1$ s and $\Delta E=1\,M_{\odot}c^{2}$ [Eq. (16)].

### v2

@@ Abstract
The history of astronomy has shown that advances in sensing methods open up new windows to the universe and often lead to unexpected discoveries. Quantum sensor networks in combination with traditional astronomical observations are emerging as a novel modality for multi-messenger astronomy. Here we develop a generic analysis framework that uses a data-driven approach to model the sensitivity of a quantum sensor network to astrophysical signals as a consequence of beyond-the-Standard Model (BSM) physics. The analysis method evaluates correlations between sensors to search for BSM signals coincident with astrophysical triggers such as black hole mergers, supernovae, or fast radio bursts. Complementary to astroparticle approaches that search for particlelike signals (e.g. WIMPs), quantum sensors are sensitive to wavelike signals from exotic quantum fields. This analysis method can be applied to networks of different types of quantum sensors, such as atomic clocks, matter-wave interferometers, and nuclear clocks, which can probe many types of interactions between BSM fields and standard model particles.
We use this analysis method to carry out the first direct search utilizing a terrestrial network of precision quantum sensors for BSM fields emitted during a black hole merger. Specifically we use the Global Network of Optical Magnetometers for Exotic physics (GNOME) to perform a search for exotic low-mass field (ELF) bursts generated in coincidence with a gravitational wave signal from a binary black hole merger (GW200311_115853) detected by LIGO/Virgo on the 11th of March 2020. The associated gravitational wave heralds the arrival of the ELF burst that interacts with the spins of fermions in the magnetometers. This enables GNOME to serve as a tool for multi-messenger astronomy. Our search found no significant events, and consequently we place the first lab-based limits on combinations of ELF production and coupling parameters.
@@ I Introduction
I.1 Overview
The Standard Model of particle physics is one of the great successes of modern science, providing a nearly perfect description of the particles that make up the stars, planets, and everyday objects around us. However, astrophysical observations have shown that this ordinary matter makes up only a small fraction of our universe Planck2018 ; Kazin . These observations strongly hint at the existence of undiscovered particles and fields beyond the Standard Model (BSM).
The last few decades have seen the development of sensors utilizing the unprecedented control of ordinary matter through the tools of quantum mechanics Ye2024 . Objects, ranging from single atoms to increasingly complex molecules gerlich2011quantum , can be placed into specific quantum states. This ability to tame the quantum nature of matter has opened up a vast toolbox enabling applications ranging from powerful quantum computers to a variety of quantum sensors with extreme precision. By clever choice of atomic species and experimental design, precise measurements of accelerations, rotations Kimball2023PRA , and electric and magnetic fields are possible. These precise quantum sensors utilizing ordinary atoms can in turn be used to search for small effects from interactions with undiscovered BSM particles and quantum fields safronova2018search . Quantum sensors are typically tabletop in size and can be deployed in networks across the world. Astrophysical signals show up as correlations across global networks allowing for the rejection of spurious backgrounds. Having a diversity of sensors can be an advantage, providing a unique fingerprint for signals from BSM physics.
Our method can search for BSM signals coincident with triggers from astronomical observations such as gravitational wave signals from black hole mergers or optical signals from fast radio bursts and even supernovae o2013interpreting . These are some of the highest energy events in the universe and provide a unique gateway for the production of BSM particles and fields in environments not accessible in Earth laboratories.
I.2 Application to ELFs
@@
Many extensions of the Standard Model posit the existence of new scalar, pseudoscalar, vector, or pseudovector fields with light ($m\ll 1\,\rm{keV/c^{2}}$) quanta antypas2022new , such as axions and axion-like particles (ALPs) PRESKILL1983127 ; ABBOTT1983133 ; DINE1983137 ; Arv10 and dark/hidden photons ackerman2009dark ; goodsell2009naturally ; nelson2011dark . In this work, we apply our general method to search for light, ultra-relativistic pseudoscalar fields triggered by an astrophysical event using the Global Network of Optical Magnetometers for Exotic physics searches (GNOME). Following Ref. Dailey2021 , we refer to these fields as exotic low-mass fields (ELFs). We define ELFs, in general, to be any BSM low-mass ($m\ll 1\,\rm{keV/c^{2}}$) bosonic field (spin-0 or spin-1). We do not assume ELFs constitute a significant fraction of dark matter and thus ELF parameters are not directly constrained by dark matter searches. Here we focus on a specific class of ELFs, namely spin-0 axion or axion-like fields that couple to fermion spins, yet the analysis method described here can be extended in straightforward ways to other classes of ELFs. [^1] Recently (during the publication process of this paper), a closely related search for ELFs was carried out using the rubidium clocks onboard satellites of the GPS sen2024multi . The search described in Ref. sen2024multi targeted a different interaction of ELFs with atoms than that probed in the present study: transient variations in fundamental constants. Analyzing GPS clock data around the GW170817 neutron star merger, the authors of Ref. sen2024multi identify an intriguing post-trigger excess in clock noise, although elevated solar electron flux during the same period complicates interpretation. This prevented them from being able to set constraints on ELF production mechanisms or interactions with atoms. ELFs are postulated to be generated during large energy astrophysical events such as supernovae Raf99 , binary black hole (BBH) or binary neutron star mergers baumann2019probing , and fast radio bursts (FRB) Tka15 . The detection of associated gravitational waves (GW) or electromagnetic (EM) signals herald the ELF signal arrival on Earth, and also indicates the GW/EM and ELF source location in the sky. By searching for signatures of the emissions from known sources shortly after a GW or EM trigger, we set limits on possible ELF production and coupling parameters associated with the observed astrophysical event.
@@
The production mechanism of ELFs is an open question arvanitaki2017black ; baryakhtar2017black ; baumann2019probing ; baryakhtar2021black ; takahashi2024self ; aurrekoetxea2024self ; eby2022probing . Black holes may be surrounded by clouds of exotic bosons, with up to 10% of the black hole mass in the clouds arvanitaki2015discovering . Theories of scalar-tensor gravity Fujii:2003pa describe cases of black holes and neutron stars being immersed in scalar fields. Such scenarios suggest that modes of these scalar fields may be excited during BBH or binary neutron star mergers franciolini2019effective . To set limits on ELF properties, we proceed with the following generic assumptions: (1) Some high-energy astrophysical event emits a spherical wave packet of the ELF $\phi(r)$ from a known location at distance $R$ from Earth. (2) The emitted wave packet carries away total energy $\Delta E$, where $\Delta E/\varepsilon_{0}\gg 1$ with $\varepsilon_{0}$ being the energy of an ELF quanta. Given the assumptions above, the ELF will have high mode occupation number and hence they are treated as classical phase-coherent waves. We place the first lab-based constraints which can be used to limit possible ELF models as well as their production mechanisms.
I.3 Outline
We present a generic analysis framework to search for correlated ELF signals in a quantum sensor network. We take a data-driven approach, using the observed background noise of the sensors to model the network sensitivity to astrophysical signals from BSM physics. We apply this approach to the specific case of searching for ELFs coincident with a GW detection with the GNOME network.
@@ I.3 Outline
This manuscript is organized as follows. In Sec. II, we recall the characteristics of an ELF pulse, its propagation, and its coupling to atomic magnetometers. In Sec. III, we describe the GNOME network. In Sec. IV, we present our two-stage analysis pipeline. First, we look for large excess power in the network at particular times and frequencies. We then perform likelihood-ratio tests using the characteristics of the individual stations to determine the significance of possible events compared to time-shifted data, where we expect no signals. We perform the same test on injected signals to calibrate the network response to a known ELF signal. Finally, in Sec. V, we present results for one search target: the BBH merger GW200311_115853 detected by LIGO/Virgo GCN_BBH .
@@ II.2 ELF detection with magnetometers
where $J^{\mu}=\bar{\psi}\gamma^{\mu}\gamma_{5}\psi$ is the axial-vector current for Standard Model fermions and $f_{l}$ and $f_{q}$ are the characteristic energy scales of the linear and quadratic interactions, respectively. Transforming to a Hamiltonian representation, we arrive at the effective spin-dependent Hamiltonians
@@
with $\bf{S}$ being the atomic or nuclear spin Afach2023 . The ELF can be treated as a so-called “pseudomagnetic field” since it causes energy shifts of Zeeman sublevels in a manner similar to the effect of a real magnetic field. [^2] In the present work we focus on the direct coupling of the ELF field $\phi$ with fermion spins cong2024spin , but interactions with other standard model particles and fields are possible, see, e.g., Refs. safronova2018search ; jacksonkimball2022 . While GNOME magnetometers afach2018characterization and especially noble-gas comagnetometers PhysRevResearch.6.013339_comag have state-of-the-art sensitivity to $\phi$-spin couplings jackson2023probing , other quantum sensors such as atomic clocks, cavities, and atom interferometers are better suited for detection of other types of $\phi$-couplings safronova2018search ; jacksonkimball2022 . One example is the search for ELF-induced transient variation of fundamental constants using the GPS atomic clock network described in Ref. sen2024multi . Another relevant example is the $\phi$-photon coupling sikivie1983experimental , which can lead to generation of a real oscillating magnetic field as opposed to a pseudomagnetic field. In this case the magnetometer need not be spin-based, and other sensors such as superconducting quantum interference devices (SQUIDs) brouwer2022proposal or induction-coil magnetometers sulai2023hunt can be used. Of course, the dependence of the measured signal on $\phi$ will depend on the details of the $\phi$-coupling and the sensor characteristics: for example, as discussed in Refs. Dailey2021 ; arakawa2025multimessenger , a scalar $\phi$-coupling that generates a signal detectable with atomic clocks does not have the directionality characteristic of the $\phi$-spin coupling considered here. To explicitly make the analogy between the Zeeman Hamiltonian ($H_{Z}=-\boldsymbol{\mu}\cdot\boldsymbol{B}$, where $\boldsymbol{\mu}=\mu\boldsymbol{S}$ is the magnetic moment and $\boldsymbol{B}$ is a magnetic field) and the Hamiltonians describing ELF-spin interactions, we introduce the corresponding pseudomagnetic fields:
@@ III GNOME Network Characteristics
GNOME comprises a global network of magnetically shielded alkali-atomic magnetometers pustelny2013global ; afach2018characterization ; Afach2021 ; Afach2023 . These are located in more than a dozen stations distributed over four continents. The sensors measure the projection of the ambient magnetic field on one sensitive axis via the variation of the Larmor spin precession of optically polarized alkali atoms in a vapor cell. Measurements are time-stamped using the Global Position System allowing for the use of coincident detection techniques to suppress unwanted background noises Afach2021 . In addition, a multi-layer mu-metal magnetic shield, present in each station, suppresses background magnetic fields by a factor of $\gtrsim 10^{6}$. Because most GNOME magnetometers use atoms whose nuclei have one unpaired proton spin, they are primarily sensitive to proton-spin couplings with the ELF. The expected ELF signal amplitudes measured by the sensors will be proportional to the relative contribution of the proton spin to the nuclear spin Afach2021 . ELFs can also couple to electron spins but the signal is attenuated by the electrons in the magnetic shields. To first order it was expected this attenuation is the same as the magnetic shielding factor kimball2016magnetic but the orbital angular momentum of the electrons in the shields can lessen this effect futurePaper . The coupling to electron spins is ignored in this manuscript.
@@
Table 1: The station name, location in longitude and latitude, orientation of the sensitive axis, and atomic species used in GNOME sensors. The rightmost column lists the estimated ratio between the effective proton spin polarization Kim15 and the Landé $g$-factor for the magnetometer, $\sigma_{p}/g$, which depends on the atomic species and the magnetometry scheme as described in Appendix B of Ref. Afach2021_domainwall . * marking denotes the stations that were active with quality data during the GW200311_115853 event. (Moved from Fribourg to Moxa in 2018.)
@@
Station       Longitude Latitude       Az Alt Species $\sigma/g_{f}$
Beersheba 34.8043° E 31.2612° N 180° 3He/87Rb/39K see footnote  [^3] The Beersheba station employs a 3He/K/Rb SERF comagnetometer similar to Ref. klinger2023optimization . The conversion between magnetic field and pseudomagnetic field for such a comagetometer, described in Ref. PhysRevResearch.6.013339_comag , is frequency dependent and therefore not described by a single number $\sigma/g_{f}$.
Beijing 116.1868° E 40.2457° N 251° 133Cs $-0.39^{+0.19}_{-0.00}$
Berkeley 122.2570° W 37.8723° N 90° 133Cs $-0.39^{+0.19}_{-0.00}$
Canberra 149.1185° E 35.2745° S 90° 87Rb $\ \ 0.70^{+0.00}_{-0.15}$
Daejeon 127.3987° E 36.3909° N 90° 133Cs $-0.39^{+0.19}_{-0.00}$
Hayward* 122.0539° W 37.6564° N 90° 87Rb $\ \ 0.70^{+0.00}_{-0.15}$
Hefei 117.2526° E 31.8429° N 90° 85Rb/87Rb $-0.38^{+0.05}_{-0.00}$
Krakow 19.9048° E 50.0289° N 45° 87Rb $\ \ 0.50^{+0.00}_{-0.11}$
Lewisburg* 76.8825° W 40.9557° N 90° 87Rb $\ \ 0.70^{+0.00}_{-0.15}$
Los Angeles* 118.4407° W 34.0705° N 270° 85Rb $\ \ 0.50^{+0.00}_{-0.07}$
Mainz 8.2354° E 49.9915° N -90° 87Rb $\ \ 0.50^{+0.00}_{-0.11}$
Moxa* 11.6147° E 50.6450° N 270° 133Cs $-0.39^{+0.19}_{-0.00}$
Oberlin* 81.7796° W 41.2950° N 276° 39K $-0.49^{+0.18}_{-0.00}$
For this work, each sensor has bandwidth of approximately 100 Hz. Magnetic field sensitivities for the sensors in each station vary from around 0.01 to 1 $\mathrm{pT/\sqrt{Hz}}$ over this band. Example characterizations are given in afach2018characterization . The response and sensitivity of each sensor is determined by calibration signals at multiple frequencies. Magnetic field measurements are recorded at a sampling rate of 512 samples per second. In addition, a 1 sample per second data glitch / data quality indicator is also written to file. This boolean variable is derived from auxiliary sensors which monitor temperature, accelerations, magnetic field outside the shielding, rotation and other user-defined variables such as laser intensity, frequency, etc. If any of those quantities vary outside an allowed range within a given second, the data-quality flag – which is normally true – is switched to false. In our analysis, we disregard all the data having data quality flag false.
@@ IV ELF Analysis Pipeline
The analysis combines methods from gravitational wave searches (EP, time-shifting) and standard statistical tools (LRT, FC construction, false positive probability) with the novel model-driven adaptively tiled spectrograms to search for ELFs with our quantum sensor network. We use data to make empirical distributions to determine the significance of candidate tiles. Note that the data analysis pipeline described here can be applied to networks of other quantum sensors searching for ELFs and also searches for ELFs composed of spin-1 particles (as opposed to the spin-0 particles assumed here) simply by appropriately modifying the function that describes the network response, e.g. Eq. (20). [^4] Another significant difference between GNOME magnetometers and other quantum sensors is their bandwidth and sample rate. For example, state-of-the-art atomic clocks generally sample data at a much lower rate while optical cavities can sample data at a much higher rate. Even for the case of GNOME, another issue is sensor dead time. For this first analysis we chose to analyze a dataset where all sensors were contributing throughout. Instead of a discrete Fourier transform, a future refinement to the algorithm could use a periodogram to analyze unevenly sampled data.
@@ V Search Target and Results
We now apply this analysis method to the BBH merger GW200311_115853 detected by LIGO/Virgo on March 11th, 2020 at 11:58:53.398 UTC PhysRevX.13.041039 ; PhysRevX.13.041039 ; GCN_BBH . The luminosity distance is estimated to be $D_{L}=1.17^{+0.28}_{-0.40}$ Gpc at right ascension 00h08m and declination -07d27m with a sky localization of 35 deg2. The total mass before the merger was $61.9^{+5.3}_{-4.2}\,M_{\odot}$ and the final mass was $59.0^{+4.8}_{-3.9}\,M_{\odot}$.
@@ V.1 Foreground analysis
Figure 9 shows the cumulative number of foreground events above a given FPP versus the mean number of expected events. The foreground and time-shifted data are largely consistent. The most significant event had an expected mean of $\mu_{tot}=0.23$ or FPP$=0.21$, which would correspond to a detection with a significance of only $1.3\sigma$. While this is not significant, we calculate the corresponding ELF masses and energy scales as an example. This tile occurred at a delay time of 15920 s and a central frequency of 95.8 Hz. Through the FC method, we obtain a 90%-confidence upper limit of $\mathcal{B}_{\text{p}}<0.08$ pT. For the linear interaction, this corresponds to a mass of $m=2.1\times 10^{-19}$ eV [Eq. (33)] and linear energy scale of $f_{l}=2.7$ GeV at $\tau_{0}=1$ s and $\Delta E=1\,M_{\odot}c^{2}$ [Eq. (15)]. For the quadratic interaction, the mass is $m=1.0\times 10^{-19}$ eV and the quadratic energy scale is $f_{q}=6.7\times 10^{-3}$ GeV at $\tau_{0}=1$ s and $\Delta E=1\,M_{\odot}c^{2}$ [Eq. (16)].